dx dt=2x-4y 

dx dt=2x+4y


1
Expert's answer
2022-06-14T13:06:37-0400
"\dfrac{dx}{dt}=2x-4y""\dfrac{dy}{dt}=2x+4y"

"Dx-2x+4y=0""Dy-2x-4y=0"

"(D-2)x+4y=0""-2x+(D-4)y=0"

Multiply first equation by "2" and second equarion by "D-2"


"2(D-2)x+8y=0""-2(D-2)x+(D-2)(D-4)y=0"

Add two equations


"-2(D-2)x+(D-2)(D-4)y"

"+(-2(D-2)x+(D-2)(D-4)y)=0"

Simplify

"D^2-6D+16y=0"

Auxiliary equation


"r^2-6r+16=0"

"r=3\pm i\sqrt{7}"

"y(t)=c_1e^{3t}\cos(\sqrt{7}t)+c_2e^{3t}\sin(\sqrt{7}t)"

Then


"\dfrac{dy}{dt}=3c_1e^{3t}\cos(\sqrt{7}t)+3c_2e^{3t}\sin(\sqrt{7}t)"

"-\sqrt{7}c_1e^{3t}\sin(\sqrt{7}t)+\sqrt{7}c_2e^{3t}\cos(\sqrt{7}t)"

Substitute


"\dfrac{dy}{dt}=2x+4y"

"x=-2y+\dfrac{1}{2}\dfrac{dy}{dt}"


"x=-2c_1e^{3t}\cos(\sqrt{7}t)-2c_2e^{3t}\sin(\sqrt{7}t)"

"+\dfrac{3}{2}c_1e^{3t}\cos(\sqrt{7}t)+\dfrac{3}{2}c_2e^{3t}\sin(\sqrt{7}t)"

"-\dfrac{\sqrt{7}}{2}c_1e^{3t}\sin(\sqrt{7}t)+\dfrac{\sqrt{7}}{2}c_2e^{3t}\cos(\sqrt{7}t)"



"x(t)=(-\dfrac{1}{2}c_1+\dfrac{\sqrt{7}}{2}c_2)e^{3t}\cos(\sqrt{7}t)"

"+(-\dfrac{1}{2}c_2-\dfrac{\sqrt{7}}{2}c_1)e^{3t}\sin(\sqrt{7}t)"

"y(t)=c_1e^{3t}\cos(\sqrt{7}t)+c_2e^{3t}\sin(\sqrt{7}t)"